Assume a 2015 Gallup Poll asked a national random sample of 482 adult women to state their current weight. Assume the mean weight in the sample was

0 0
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 164

Population standard deviation = \sigma = 36

Sample size = n = 482

Z\alpha/2 = 1.645

Margin of error = E = Z\alpha/2* (\sigma /\sqrtn)

= 1.645 * (36 / \sqrt 482)

= 2.70

At 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

164 - 2.70 < \mu < 160 + 2.70

161.3 < \mu < 162.70

Lower value = 161.3

Upper value = 162.70